I am trying to figure out this equation.
3x/4+1=10

Answer by checkley77(12844)

3x/4+1=10
3x/4=10-1
3x/4=9
3x=4*9
3x=36
x=36/3
x=12 answer.
Proof:
3*12/4+1=10
36/4+1=10
9+1=10
10=10